# Logarithmic regret in the ergodic Avellaneda–Stoikov market making model

### v1

@@ Abstract.
We analyse the regret arising from learning the price sensitivity parameter $\kappa$ of liquidity takers in the ergodic version of the Avellaneda–Stoikov market making model. We show that a learning algorithm based on a regularised maximum-likelihood estimator for the parameter achieves the regret upper bound of order $\ln^{2}T$ in expectation. To obtain the result we need two key ingredients. The first are tight upper bounds on the derivative of the ergodic constant in the Hamilton–Jacobi–Bellman (HJB) equation with respect to $\kappa$. The second is the learning rate of the maximum-likelihood estimator which is obtained from concentration inequalities for Bernoulli signals. Numerical experiment confirms the convergence and the robustness of the proposed algorithm.
@@ 1. Introduction
where $(S_{t})_{t\geq 0}$ is the exogenous mid-price process, $(Q_{t}^{\delta^{\pm}})_{t\geq 0}$ is the market maker’s inventory and $(X^{\delta^{\pm}}_{t})_{t\geq 0}$ is the market maker’s cash balance. The inventory and cash processes are driven by $N_{t}^{\delta,\pm}$, two independent Poisson jump processes with intensities $\lambda^{\pm}e^{-\kappa\delta^{\pm}}$. The market maker wishes to maximize the long-run average reward which sums the earnings and changes to mark-to-market value of their holdings of the risky asset but is subject to a quadratic inventory penalty expressing their risk aversion:
@@
The model parameters are: liquidity takers orders’ arrival rates $\lambda^{\pm}$, the price sensitivity of the liquidity takers $\kappa$, the mid price volatility $\sigma$ (which actually plays no role as the mid price process is a martingale) and the risk aversion $\phi$. The market maker chooses their risk aversion and thus it’s not a parameter that they would need to learn. The liquidity takers orders’ arrival rates $\lambda^{\pm}$ can be observed and learned offline (without participating in the market) since, in the framework of the model where our market maker is assumed to provide a relatively small fraction of the overall liquidity, it is unlikely that the presence of their volume in the market would impact the rate of liquidity taking. This leaves $\kappa$ and this is the key parameter. It is not reasonable to assume that this could be learned offline. Exchanges provide visibility of the order book and some information on executed trades but often not enough to allow good offline estimate for $\kappa$. Furthermore other market makers will react to the presence of the additional volume placed by our market maker at the distance $\delta^{\pm}$ thus potentially rendering any offline estimate of $\kappa$ invalid. Indeed, some may choose to place their volume at better price (they want to trade) than the spread given by our market maker while others may wish to place the volume at worse price (they may think our market maker knows something about the price they don’t).
The key parameter to learn online (i.e. while participating in the market) is thus $\kappa$. At each time $t\geq 0$ the market maker will have their estimate of the parameter denoted $\kappa_{t}$ while the true, unknown, value is $\kappa^{\ast}$. They can solve the ergodic control problem and obtain the strategy which would be optimal if $\kappa_{t}$ would be the true parameter. Let us denote this by strategy by $\psi^{\kappa_{t},\pm}$.
@@
(1) $\mathcal{R}(T)=\gamma(\kappa^{\star})T-\mathbb{E}_{q,x,S}\bigg{[}\int_{0}^{T}\mathrm{d}\big{(}X^{\psi^{\kappa_{t},\pm}}_{t}+S_{t}Q^{\psi^{\kappa_{t},\pm}}_{t}\big{)}-\phi\int_{0}^{T}(Q^{\psi^{\kappa_{t},\pm}}_{t})^{2}\,\mathrm{d}t\bigg{]}\,.$
@@
If the market maker would use a fixed $\kappa\neq\kappa^{\ast}$ then their expected regret would be roughly $(\gamma(\kappa^{\ast})-\gamma(\kappa))T$ i.e. linear. Any algorithm which achieves sub-linear regret is learning. We construct a regularised maximum-likelihood estimator, see (25), and employ it in Algorithm 1, to achieve the expected regret upper bound of order $\ln^{2}T$. See Theorem 19.
@@
There are three main ingredients which allow us to obtain this result. First, we prove existence of and convergence to an invariant measure in the ergodic Avellaneda–Stoikov market making model. While the well-posedness of the ergodic problem follows mostly from the analysis carried out in Guéant and Manziuk [27] the result on existence of and convergence to the invariant measure is new and relies on newly established explicit solution the the ergodic HJB corresponding to our problem.
Second, we obtain bounds on the second derivative of the average earnings per unit time (the ergodic constant) with respect to the parameter $\kappa$ which has to be learned.
@@ 2. Main results
The filtration is $\mathbb{F}:=(\mathcal{F}_{t})_{t\geq 0}$, where $\mathcal{F}_{t}=\sigma(W_{r},r\leq t)\vee\sigma(M_{r}^{\pm},r\leq t)\vee\sigma(U^{+}_{M^{+}_{r}},r\leq t)\vee\sigma(U^{-}_{M^{-}_{r}},r\leq t)$. Let $\underline{q}\in\mathbb{Z}^{-}$ and $\bar{q}\in\mathbb{Z}^{+}$ denote the market maker’s inventory limits. The market maker will stop posting buy/sell orders when their inventory is at $\bar{q}$ and at $\underline{q}$ respectively. At other times their strategy is to post at a distance $\delta^{\pm}\in\mathbb{R}$ from the mid-price $S_{t}$. Clearly, the strategy $\delta^{\pm}$ must be adapted to the filtration $\mathbb{F}$. Let $(N_{t}^{\delta,\pm})_{t\geq 0}$ be the controlled counting processes for the agent’s filled buy/sell orders, i.e.
$N_{t}^{\delta,\pm}=N_{t-}^{\delta,\pm}+(M^{\pm}_{t}-M^{\pm}_{t-})\bm{1}_{\left\{U^{\pm}_{M^{\pm}_{t}}\geq e^{-\kappa^{\pm}\delta_{t-}^{\pm}}\right\}}\,.$
@@ 2.1. The ergodic market making model
In this section we will formulate the ergodic control problem, state key results connecting the control formulation with the ergodic HJB equation, provide explicit solution for the ergodic HJB and formulae for the Markovian ergodic optimal controls, state equilibrium properties of the system and state a result on convergence to equilibrium.
@@
The market maker aims to maximise a long-run average reward of the accumulated PnL with a running inventory penalty. Let $J(x,S,q;\delta^{\pm})$ be the ergodic reward functional given by
@@
Key results for the discounted finite time and infinite-time problems
@@
(10) $\begin{split}H(q,\delta^{\pm},\bm{p})&=\lambda^{+}e^{-\kappa^{+}\delta^{+}}(p_{1}+\delta^{+})\bm{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{-}\delta^{-}}(p_{2}+\delta^{-})\bm{1}_{q<\bar{q}}-\phi q^{2}\,,\\ H(q,\bm{p})&=\sup_{\delta^{\pm}\in\mathbb{R}^{2}}H(q,\delta^{\pm},\bm{p})\,.\end{split}$
@@
The ergodic HJB and its connection the ergodic control problem
In this section, we analyse the ergodic control problem (9) by considering the asymptotic behavior of $T\rightarrow+\,\infty$ in the finite-time-horizon model (12) with $r=0$ and $r\rightarrow 0$ in the discounted infinite-time-horizon model (14). We prove that $\lim_{r\rightarrow 0}rv_{r}(q)$ is equal to the ergodic constant $\gamma$ in (9). Then explicit solutions to the ergodic control problem are derived.
@@
So far we’ve established the connection between the ergodic constant $\gamma$ and the solution of the ergodic HJB equation (16). Next, we are interested in how this constant depends on the model parameter $\kappa$. This is best seen from an explicit formulation for $\gamma=\gamma(\kappa)$ given in the following theorem.
@@
$\gamma=\gamma(\kappa)=\frac{\lambda_{max}(\kappa)}{\kappa}\,.$
@@
The optimal feedback (Markov) control for the ergodic control problem $\psi^{\pm}$ is uniquely given by
@@
Obviously, $\psi^{\pm}$ given by (18) depends on the model parameters, in particular on $\kappa$. We will denote the optimal feedback control for the ergodic problem with the parameter $\kappa$ as $\psi^{\pm,\kappa}$.
@@
$\bm{C}=\begin{bmatrix}-\kappa(\phi\bar{q}^{2}+\gamma)&\lambda^{+}e^{-1}&0&&&...&\\ \lambda^{-}e^{-1}&-\kappa\big{(}\phi(\bar{q}-1)^{2}+\gamma\big{)}&\lambda^{+}e^{-1}&&&...&\\ &&&...&\\ &...&&&\lambda^{-}e^{-1}&-\kappa\big{(}\phi(\underline{q}+1)^{2}+\gamma\big{)}&\lambda^{+}e^{-1}\\ &...&&&0&\lambda^{-}e^{-1}&-\kappa(\phi\underline{q}^{2}+\gamma)\end{bmatrix}$
@@
i.e. $\bm{\hat{\omega}}$ is the non-trivial solution to the homogeneous equation with coefficient matrix $\bm{C}$.
Notice that once we’ve obtained $\bm{\hat{\omega}}$ by solving (19) we have an explicit formula for the optimal ergodic control $\psi^{\pm}$ uniquely given by (18).
Convergence to equilibrium
We’ve derived the explicit solutions to the ergodic control problem (9). Now we proceed to analyse the ergodicity of the state $(Q_{t})_{t\geq 0}\in\Omega^{Q}$, i.e. how fast the state $(Q_{t})_{t\geq 0}$ converges to the equilibrium distribution. This will be essential in the regret analysis. Clearly, the state space $\Omega^{Q}$ is discrete, hence we can utilise the classical discrete state-space Markov chain theory. Let us first define the equilibrium under the Markov control $\delta^{\pm}$ for the market making system as follows.
Definition 10 (Equilibrium).
The following lemma is proved in Appendix A.10.
Lemma 11.
There exists an equilibrium distribution under the ergodic optimal control $\psi^{\pm}$ given by (18).
As we show in Appendix A.10, the dynamics of the inventory (4) can be considered as a discrete Markov chain with the transition probability matrix $\bm{P}$ defined in (41). For any $i\in\{0,1,...,\bar{q}-\underline{q}\}$, we have $\bm{P}_{ii}>0$. Hence, by definition, $\bm{P}$ is aperiodic. Moreover, as $\bm{P}$ is a tridiagonal matrix, for any $N\geq(\bar{q}-\underline{q}+1)$, $\bm{P}^{N}$ has all positive entries, i.e. $\forall i,j\in\{0,1,...,\bar{q}-\underline{q}\}$, $\bm{P}^{N}_{ij}>0$. Hence $\bm{P}$ is irreducible. Therefore, the convergence rate for any distribution under $\bm{P}$ to the equilibrium distribution satisfies the so-called Convergence Theorem [21].
Lemma 12 (Convergence Theorem).
Let $\pi_{t}$ be the probability distribution of the random variables $Q_{t}$ that follows the controlled dynamics (4) with an initial state $Q_{0}\sim\pi_{0}$ and the ergodic optimal control (18). Let $\pi$ be the equilibrium distribution established by Lemma 11. Then, there exists constants $C>0$ and $10<\alpha<$ 1 such that
$\left\|\pi_{t}-\pi\right\|_{TV}\leq C\alpha^{t},\quad\forall t\geq 0\,.$
@@ 2.2. Learning and regret
In this section, we consider the parameter learning problem of the market making model in the ergodic setting, where the price sensitivity of market participants parameter is unknown to the market maker. We assume that the parameter is equal on the bid/ask side $\kappa^{\star}=\kappa^{\star,\pm}$ and $k^{\star}\in[\underline{K},\bar{K}]$ for some $\underline{K},\bar{K}\in\mathbb{R}^{+}$ with $\underline{K}\leq\bar{K}$. The market maker does not observe $\kappa^{\star}$, but has a prior knowledge that $\kappa^{\star}\in[\underline{K},\bar{K}]$. At each time $t>0$, the market maker generates the estimate of the parameter denoted $\kappa_{t}$ in the compact set $[\underline{K},\bar{K}]$ from the learning algorithm, see Algorithm 1 for more details. By assuming this is the true parameter they can solve the ergodic control problem and obtain the policy $\psi^{\kappa_{t}}$ given by (18) which would be optimal if $\kappa_{t}$ indeed was the true parameter.
@@
In a general RL problem the agent aims to learn from data, e.g. states, actions and rewards, a policy that optimises the reward, see [10, 38, 24]. In this learning problem, we have derived the global optimal policy $\psi^{\pm}$ (see Section 2.1). The global optimal policy is attainable if the true $\kappa^{\star}$ is known. Therefore, it is sufficient to define a learning algorithm to generate the parameter $\kappa$.
@@
Let $(\Omega^{\star},\mathcal{F}^{\star},\mathbb{P}^{\star})$ be defined as
$(\Omega^{\star},\mathcal{F}^{\star},\mathbb{P}^{\star})=\big{(}\Omega^{M}\times\Omega^{U},\mathcal{F}^{M}\otimes\mathcal{F}^{U},\mathbb{P}^{M}\otimes\mathbb{P}^{U}\big{)}\,,$
see details in (3), $\mathcal{N}$ be the $\sigma-$algebra generated by $\mathbb{P}^{\star}-$null sets, and the continuous-time learning algorithm $\Psi=(\Psi_{t})_{t}$ be some function $\Psi:\Omega^{\star}\times\mathbb{R}^{+}\mapsto[\underline{K},\bar{K}]$. We say that $\Psi=(\Psi_{t})_{t}$ is an admissible learning algorithm if $\Psi$ is $\big{(}\mathcal{G}^{\Psi}_{t^{-}}\otimes\mathcal{B}(\mathbb{R}^{+})\big{)}/\mathcal{B}([\underline{K},\bar{K}])$ measurable, with the $\sigma-$algebra $\mathcal{G}^{\Psi}=(\mathcal{G}^{\Psi}_{t})_{t}$ defined as $\mathcal{G}^{\Psi}_{t}:=\sigma\big{\{}M_{s}^{\Psi;\kappa^{\star},\pm}\big{|}0<s\leq t\big{\}}\vee\sigma\big{\{}U_{M^{\Psi;\kappa^{\star},\pm}_{s}}^{\Psi;\kappa^{\star},\pm}\big{|}0<s\leq t\big{\}}\vee\mathcal{N}$.
@@
$\mathcal{G}_{t}^{\Psi}$ in Definition 14 describes the available and useful information for the agent to estimate $\kappa$ up to time $t$. Moreover, it is not hard to see [31, 40] that the learning algorithm $\kappa_{t}$ generated by a maximum likelihood estimator is $\mathcal{G}^{\Psi}_{t^{-}}-$measurable.
@@
(20) $\mathcal{R}^{\Psi}(T)=\gamma(\kappa^{\star})T-\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}},\psi^{\kappa_{t}};\kappa^{\star})\,\mathrm{d}t\Big{]}\,,$
where $\gamma(\kappa^{\star})$ is the optimal long-run average reward under the parameter $\kappa^{\star}$, $f$ is the running reward function given by
(21) $f(q,\delta^{\pm};\kappa^{\star})=\lambda^{+}\delta^{+}e^{-\kappa^{\star}\delta^{+}}+\lambda^{-}\delta^{-}e^{-\kappa^{\star}\delta^{-}}-\phi q^{2}\,,$
and $Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}}$ is the inventory process governed by $\kappa^{\star}$ but with the control $\psi^{\kappa_{t}}$, i.e.
(22) $\begin{split}dQ_{t}^{\psi^{\kappa};\kappa^{\star}}&=dN_{t}^{\psi^{\kappa};\kappa^{\star},-}-dN_{t}^{\psi^{\kappa};\kappa^{\star},+}\\ &=\big{(}\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,-}}-\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,+}}\big{)}dt+d\tilde{N}_{t}^{\psi^{\kappa};\kappa^{\star},-}-d\tilde{N}_{t}^{\psi^{\kappa};\kappa^{\star},+}\,,\end{split}$
with $N_{t}^{\psi^{\kappa};\kappa^{\star},\pm}$ the controlled counting processes for the market maker’s filled buy/sell orders and $\tilde{N}_{t}^{\psi^{\kappa};\kappa^{\star},\pm}$ the corresponding compensated Poisson processes.
@@
(23) ^$\begin{split}\widehat{\mathcal{R}}^{\Psi}(T)&=J(\psi^{\kappa^{\star}};\kappa^{\star})-J(\psi^{\kappa_{t}};\kappa^{\star})\\ &=\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(Q_{t}^{\psi^{\kappa^{\star}};\kappa^{\star}},\psi^{\kappa^{\star}};\kappa^{\star})\,\mathrm{d}t\Big{]}-\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}},\psi^{\kappa_{t}};\kappa^{\star})\,\mathrm{d}t\Big{]}\,.\end{split}$
@@
(24) $\left|\gamma(\kappa^{\star})T-\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(Q_{t}^{\psi^{\kappa^{\star}};\kappa^{\star}},\psi^{\kappa^{\star}};\kappa^{\star})\,\mathrm{d}t\Big{]}\right|\leq C\,,$

### v2

@@ Abstract.
We analyse the regret arising from learning the price sensitivity parameter $\kappa$ of liquidity takers in the ergodic version of the Avellaneda–Stoikov market making model. We show that a learning algorithm based on a maximum-likelihood estimator for the parameter achieves the regret upper bound of order $\ln^{2}T$ in expectation. To obtain the result we need two key ingredients. The first is the twice differentiability of the ergodic constant under the misspecified parameter in the Hamilton–Jacobi–Bellman (HJB) equation with respect to $\kappa$, which leads to a second–order performance gap. The second is the learning rate of the regularised maximum-likelihood estimator which is obtained from concentration inequalities for Bernoulli signals. Numerical experiments confirm the convergence and the robustness of the proposed algorithm.
@@ 1. Introduction
The model was originally formulated in a finite-time-horizon setting, where the market maker’s objective is to maximise the expected profit over a fixed time period. In this paper we re-formulate the model in an ergodic setting. To formulate a learning algorithm and its regret in the finite-time-horizon model we would have considered the episodic setting. That is, the market maker runs with a fixed $\kappa$ until the time $T$ and then liquidates their inventory, updates their estimate of $\kappa$ and starts again. This feels unnatural as liquidating the entire inventory at $T$ with a market order would be costly and not the behaviour one would expect. It seems more realistic to assume that the market maker is continuously learning the parameter $\kappa$ and updating their strategy based on the new information while managing their inventory risk according to their risk appetite expressed via a quadratic penalty on the inventory and the inventory bounds. The ergodic formulation allows us to capture learning and regret in this more natural setting.
@@
where $(S_{t})_{t\geq 0}$ is the exogenous mid-price process, $(Q_{t}^{\delta^{\pm}})_{t\geq 0}$ is the market maker’s inventory and $(X^{\delta^{\pm}}_{t})_{t\geq 0}$ is the market maker’s cash balance. The inventory and cash processes are driven by $N_{t}^{\delta,\pm}$, two independent Poisson jump processes with intensities $\lambda^{\pm}e^{-\kappa\delta^{\pm}}$. The market maker wishes to maximise the long-run average reward which sums the earnings and changes to mark-to-market value of their holdings of the risky asset but is subject to a quadratic inventory penalty expressing their risk aversion:
@@
The model parameters are: liquidity takers orders’ arrival rates $\lambda^{\pm}$, the price sensitivity of the liquidity takers $\kappa$, the mid price volatility $\sigma$ (which actually plays no role as the mid price process is a martingale) and the risk aversion $\phi$. The market maker chooses their risk aversion and thus it’s not a parameter that they would need to learn. The liquidity takers orders’ arrival rates $\lambda^{\pm}$ can be observed and learned offline (without participating in the market) since, in the framework of the model where our market maker is assumed to provide a relatively small fraction of the overall liquidity, it is unlikely that the presence of their volume in the market would impact the rate of liquidity taking. This leaves $\kappa$ and this is the key parameter. Although exchanges may provide market participants with visibility of the order book and message-level trades execution data, allowing them to estimate $\kappa$ without direct participation, this is insufficient for an accurate estimate for $\kappa$. A key challenge is that other market makers will react to the presence of the additional volume placed by our market maker at the distance $\delta^{\pm}$ thus potentially rendering any offline estimate of $\kappa$ inaccurate. Indeed, some liquidity providers may choose to place their volume at better price (they want to trade) than the spread given by our market maker while others may wish to place the volume at worse price (they may think our market maker knows something about the price they don’t). The offline estimate of $\kappa$ can of course be used as the initial value in the learning algorithm.
The key parameter to learn online (i.e. while participating in the market) is thus $\kappa$. At each time $t\geq 0$ the market maker will have their estimate of the parameter denoted $\kappa_{t}$ while the true, unknown, value is $\kappa^{\ast}$. They can solve the ergodic control problem and obtain the strategy which would be optimal if $\kappa_{t}$ would be the true parameter. Let us denote this strategy by $\psi^{\kappa_{t},\pm}$.
@@
(1) $\mathcal{R}(T)=\gamma(\kappa^{\ast})T-\mathbb{E}_{q,x,S}\bigg{[}\int_{0}^{T}\mathrm{d}\big{(}X^{\psi^{\kappa_{t},\pm}}_{t}+S_{t}Q^{\psi^{\kappa_{t},\pm}}_{t}\big{)}-\phi\int_{0}^{T}(Q^{\psi^{\kappa_{t},\pm}}_{t})^{2}\,\mathrm{d}t\bigg{]}\,.$
@@
If the market maker would use a fixed $\kappa\neq\kappa^{\ast}$ then their expected regret would be roughly $(\gamma(\kappa^{\ast})-\gamma(\kappa;\kappa^{\ast}))T$, i.e. linear. Any algorithm which achieves sub-linear regret is learning. We construct a regularised maximum-likelihood estimator, see (28) and Algorithm 1, to achieve the expected regret upper bound of order $\ln^{2}T$. See Theorem 17.
@@
There are three main ingredients which allow us to obtain this result. First, we prove existence of and convergence to an invariant measure in the ergodic Avellaneda–Stoikov market making model. While the well-posedness of the ergodic problem follows mostly from the analysis carried out in Guéant and Manziuk [26] the result on existence of and convergence to the invariant measure is new and relies on newly established explicit solution to the ergodic HJB corresponding to our problem.
Second, we obtain bounds on the second–order derivative of the average earnings per unit time (i.e. the ergodic constant under a misspecified $\kappa$) with respect to the parameter $\kappa$ which has to be learned. This leads to a second-order performance gap in the regret analysis, which is crucial.
@@ 2. Main results
The filtration is $\mathbb{F}:=(\mathcal{F}_{t})_{t\geq 0}$, where $\mathcal{F}_{t}=\sigma(W_{r},r\leq t)\vee\sigma(M_{r}^{\pm},r\leq t)\vee\sigma(U^{+}_{M^{+}_{r}},r\leq t)\vee\sigma(U^{-}_{M^{-}_{r}},r\leq t)$. Let $\underline{q}\in\mathbb{Z}^{-}$ and $\bar{q}\in\mathbb{Z}^{+}$ denote the market maker’s inventory limits. The market maker will stop posting buy/sell orders when their inventory is at $\bar{q}$ and at $\underline{q}$ respectively. At other times their strategy is to post at a distance $\delta^{\pm}\in\mathbb{R}$ from the mid-price $S_{t}$. The reason for imposing inventory boundaries is that they reduce an infinite state-space control problem into a finite one, making it computationally tractable by leading to a matrix representation for an explicit solution. Clearly, the strategy $\delta^{\pm}$ must be adapted to the filtration $\mathbb{F}$. Let $(N_{t}^{\delta,\pm})_{t\geq 0}$ be the controlled counting processes for the agent’s filled buy/sell orders, i.e.
$N_{t}^{\delta,\pm}=N_{t-}^{\delta,\pm}+(M^{\pm}_{t}-M^{\pm}_{t-})\mathds{1}_{\left\{U^{\pm}_{M^{\pm}_{t}}\geq e^{-\kappa^{\pm}\delta_{t-}^{\pm}}\right\}}\,.$
@@ 2.1. The ergodic market making model
In this section we will formulate the ergodic control problem, state key results connecting the control formulation with the ergodic HJB equation, provide explicit solution for the ergodic HJB and formulae for the Markovian ergodic optimal controls.
@@
The market maker aims to maximise a long-run average reward of the accumulated PnL with a running inventory penalty. The quadratic penalty on running inventory plays a crucial role by providing a continuous incentive to steadily drive the inventory level toward zero. This is important in any volatile market, where the market maker seeks to minimise directional exposure to adverse price movements.
Let $J(x,S,q;\delta^{\pm})$ be the ergodic reward functional given by
@@
Key results for the discounted finite-time and infinite-time problems
@@
(10) $\begin{split}H(q,\delta^{\pm},\bm{p})&=\lambda^{+}e^{-\kappa^{+}\delta^{+}}(p_{1}+\delta^{+})\mathds{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{-}\delta^{-}}(p_{2}+\delta^{-})\mathds{1}_{q<\bar{q}}-\phi q^{2}\,,\\ H(q,\bm{p})&=\sup_{\delta^{\pm}\in\mathbb{R}^{2}}H(q,\delta^{\pm},\bm{p})\,.\end{split}$
@@
The ergodic HJB and its connection to the ergodic control problem
In this section, we analyse the ergodic control problem (9) by considering the asymptotic behaviour of $T\rightarrow+\,\infty$ in the finite-time-horizon model (12) with $r=0$ and $r\rightarrow 0$ in the discounted infinite-time-horizon model (14). We prove that $\lim_{r\rightarrow 0}rv_{r}(q)$ is equal to the ergodic constant $\gamma$ in (9). Then explicit solutions to the ergodic control problem are derived.
@@
So far we’ve established the connection between the ergodic constant $\gamma$ and the solution to the ergodic HJB equation (16). Next, we are interested in how this constant depends on the model parameter $\kappa$. This is best seen from an explicit formulation for $\gamma=\gamma(\kappa)$ given in the following theorem.
@@
(18) $\gamma=\gamma(\kappa)=\frac{\lambda_{max}(\kappa)}{\kappa}\,.$
@@
The optimal feedback (Markov) control for the ergodic control problem $\psi=(\psi^{+},\psi^{-})$ is uniquely given by
@@
Obviously, $\psi$ given by (19) depends on the model parameters, in particular on $\kappa$. We will denote the optimal feedback control for the ergodic problem with the parameter $\kappa$ as $\psi^{\kappa}$.
@@
$\bm{C}=\begin{bmatrix}-\kappa(\phi\bar{q}^{2}+\gamma)&\lambda^{+}e^{-1}&0&&&...&\\ \lambda^{-}e^{-1}&-\kappa\big{(}\phi(\bar{q}-1)^{2}+\gamma\big{)}&\lambda^{+}e^{-1}&&&...&\\ &&&...&\\ &...&&&\lambda^{-}e^{-1}&-\kappa\big{(}\phi(\underline{q}+1)^{2}+\gamma\big{)}&\lambda^{+}e^{-1}\\ &...&&&0&\lambda^{-}e^{-1}&-\kappa(\phi\underline{q}^{2}+\gamma)\end{bmatrix}\,.$
@@
i.e. $\bm{\hat{\omega}}$ is the non-trivial solution to the homogeneous equation with coefficient $\bm{C}$. Moreover, $\bm{\hat{\omega}}$ can be chosen to be positive, and it is unique up to a scalar multiple.
Notice that once we’ve obtained $\bm{\hat{\omega}}$ by solving (20) we have an explicit formula for the optimal ergodic control $\psi$ uniquely given by (19).
@@ 2.2. Learning and regret
In this section, we consider the parameter learning problem of the market making model in the ergodic setting, where the price sensitivity of the liquidity takers is unknown to the market maker. We assume that the parameter is equal on the bid/ask side $\kappa^{\ast}=\kappa^{\ast,\pm}\in\mathbb{R^{+}}$. The market maker does not observe $\kappa^{\ast}$, but works with the prior assumption that $\kappa^{\ast}$ must be in $[\underline{K},\bar{K}]$ with $0<\underline{K}<\bar{K}$. At each time $t>0$, the market maker generates the estimate of the parameter denoted $\kappa_{t}$ from the regularised maximum–likelihood estimator, see Algorithm 1 for more details. Using $\kappa_{t}$ they can solve the ergodic control problem and obtain the policy $\psi^{\kappa_{t}}$ given by (19).
@@
In a general RL problem the agent aims to learn from data, e.g. states, actions and rewards, a policy that optimises the reward, see [11, 39, 23]. In this learning problem, we have derived the global optimal policy $\psi$ (see Section 2.1). The global optimal policy is attainable if the true $\kappa^{\ast}$ is known. Therefore, it is sufficient to define a learning algorithm to generate the parameter $\kappa$.
@@
Let $(\Omega^{\ast},\mathcal{F}^{\ast},\mathbb{P}^{\ast})$ be defined as
$(\Omega^{\ast},\mathcal{F}^{\ast},\mathbb{P}^{\ast})=\big{(}\Omega^{M}\times\Omega^{U},\mathcal{F}^{M}\otimes\mathcal{F}^{U},\mathbb{P}^{M}\otimes\mathbb{P}^{U}\big{)}\,,$
see details in (3), $\mathcal{N}$ be the $\sigma$-algebra generated by $\mathbb{P}^{\ast}-$null sets, and the continuous-time learning algorithm $\Psi=(\Psi_{t})_{t}$ be some function $\Psi:\Omega^{\ast}\times\mathbb{R}^{+}\rightarrow[\underline{K},\bar{K}]$. We say that $\Psi=(\Psi_{t})_{t}$ is an admissible learning algorithm if $\Psi$ is $\big{(}\mathcal{G}^{\Psi}_{t^{-}}\otimes\mathcal{B}(\mathbb{R}^{+})\big{)}/\mathcal{B}([\underline{K},\bar{K}])$ measurable with the $\sigma$-algebra $\mathcal{G}^{\Psi}=(\mathcal{G}^{\Psi}_{t})_{t}$ defined as $\mathcal{G}^{\Psi}_{t}:=\sigma\big{\{}M_{s}^{\Psi;\kappa^{\ast},\pm}\big{|}0<s\leq t\big{\}}\vee\sigma\big{\{}U_{M^{\Psi;\kappa^{\ast},\pm}_{s}}^{\Psi;\kappa^{\ast},\pm}\big{|}0<s\leq t\big{\}}\vee\mathcal{N}$.
@@
$\mathcal{G}_{t}^{\Psi}$ in Definition 11 describes the available and useful information for the agent to estimate $\kappa$ up to time $t$. Moreover, it is not hard to see [31, 41] that the learning algorithm $\kappa_{t}$ generated by a maximum likelihood estimator is $\mathcal{G}^{\Psi}_{t^{-}}$ measurable.
@@
(21) $\mathcal{R}^{\Psi}(T)=\gamma(\kappa^{\ast})T-\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}},\psi^{\kappa_{t}};\kappa^{\ast})\,\mathrm{d}t\Big{]}\,,$
where $\gamma(\kappa^{\ast})$ is the optimal long-run average reward under the parameter $\kappa^{\ast}$, $f$ is the running reward function given by
(22) $f(q,\delta^{\pm};\kappa^{\ast})=\lambda^{+}\delta^{+}e^{-\kappa^{\ast}\delta^{+}}+\lambda^{-}\delta^{-}e^{-\kappa^{\ast}\delta^{-}}-\phi q^{2}\,,$
and $Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}}$ is the inventory process governed by $\kappa^{\ast}$ but with the control $\psi^{\kappa_{t}}$, i.e.
(23) $\begin{split}dQ_{t}^{\psi^{\kappa};\kappa^{\ast}}&=dN_{t}^{\psi^{\kappa};\kappa^{\ast},-}-dN_{t}^{\psi^{\kappa};\kappa^{\ast},+}\\ &=\big{(}\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,-}}-\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,+}}\big{)}dt+d\tilde{N}_{t}^{\psi^{\kappa};\kappa^{\ast},-}-d\tilde{N}_{t}^{\psi^{\kappa};\kappa^{\ast},+}\,,\end{split}$
with $N_{t}^{\psi^{\kappa};\kappa^{\ast},\pm}$ the controlled counting processes for the market maker’s filled buy/sell orders and $\tilde{N}_{t}^{\psi^{\kappa};\kappa^{\ast},\pm}$ the corresponding compensated Poisson processes.
@@
(24) ^$\begin{split}\widehat{\mathcal{R}}^{\Psi}(T)&=J(\psi^{\kappa^{\ast}};\kappa^{\ast})-J(\psi^{\kappa_{t}};\kappa^{\ast})\\ &=\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(Q_{t}^{\psi^{\kappa^{\ast}};\kappa^{\ast}},\psi^{\kappa^{\ast}};\kappa^{\ast})\,\mathrm{d}t\Big{]}-\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}},\psi^{\kappa_{t}};\kappa^{\ast})\,\mathrm{d}t\Big{]}\,.\end{split}$
@@
(25) $\left|\gamma(\kappa^{\ast})T-\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(Q_{t}^{\psi^{\kappa^{\ast}};\kappa^{\ast}},\psi^{\kappa^{\ast}};\kappa^{\ast})\,\mathrm{d}t\Big{]}\right|\leq C\,,$